OpenStudy (anonymous):

If g(x) = .5x + 3, find the value of g(0) + 2. A. 5 B. 4 C. 5.5 D. 7

OpenStudy (anonymous):

\[g(0)=.5\times 0+3=3\]so \[g(0)+2=3+2=5\]